Rumolt Motors has 29 million shares outstanding with a share price of $52 per share. In addition, Rumolt has issued bonds with a total current market value of $1,212 million. Suppose Rumolts equity cost of capital is 10%, and its debt cost of capital is 6%. a. What is Rumolts pre-tax WACC? b. If Rumolts corporate tax rate is 35%, what is its after-tax WACC?
Answer #1

Total value of equity=(29million shares*$52)=$1508million

Hence total value=(1508+1212)=$2720million

a.Pretax WACC=Respective costs*Respective weights

=(1508/2720*10)+(1212/2720*6)

=8.22%(Approx).

b.After-tax cost of debt=6(1-tax rate)

=6(1-0.35)=3.9%

WACC=Respective costs*Respective weights

=(1508/2720*10)+(1212/2720*3.9)

=7.28%(Approx).